# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 7 Conclusion and Future Work
We have given bounds for $\lambda_{2}^{*}$ (and often, $\gamma^{(1)}$) for a large number of graph classes, as well as a new relationship between $\lambda_{2}^{*}$ and $\gamma^{(1)}$. There are a number of natural questions left open, besides those previously discussed. Some of the most far-reaching questions are as follows:
  • The new relationship between $\lambda_{2}^{*}(G)$ and $\gamma^{(1)}(G)$ incorporating $\alpha(G)$, stated in Theorem 5.3, is in some ways similar to an approximation algorithm of Feige, Hajiaghayi, and Lee [FHL05, FHL08]. In particular, it is natural to ask if vertex expansion may be $\operatorname{\mathcal{O}}(\alpha(G))$-approximated in almost-linear time by strengthening the main result of Lau, Tung, and Wang [LTW24], which gives an almost-linear time $\operatorname{\mathcal{O}}(\sqrt{\log n})$-approximation. In particular, both results build on the metric rounding method Arora, Rao, and Vazirani [ARV08, ARV09].
  • Similarly, it would be interesting to determine if Theorem 5.3 could be made directly constructive, ideally with a simple algorithm. Specifically, this would remove the algorithmic requirements of bounds on $\alpha(G)$. Such a construction would not necessarily need to compute $\alpha(G)$.
  • While the $\log\Delta$ factor in Theorem 5.3 is likely essentially necessary in general graphs [LRV13], we have shown that at least one other possible factor may be used ($\alpha(G)$) that is stronger in some cases. It would be interesting to know if other factors may be used in this dimension-reduction bound, such as the logarithm of average degree, or the logarithm of the square root of the average squared degree. It would also be interesting to know if there is a natural way to unify all these bounds in some manner, particularly if such a unification could result in one simple algorithm.
  • Can these methods be leveraged to efficiently obtain near-optimal balanced separators? The methods of Orecchia and Vishnoi [OV11], and Orecchia, Sachdeva, and Vishnoi [OSV12] seem to be promosing avenues for this possibility.
In addition, there is the practical question of whether or not reweighted spectral partitioning produces better results than other techniques in real-world settings. The methods for reweighted spectral partitioning are still quite new, and has primarily been studied primarily from a theoretical perspective. Hence, this significantly differs from the history of spectral partitioning, which was shown to have practical significance before being shown to have algorithmic guarantees.
Experimentally testing the practicality of the method would first require a complete implementation, so we outline the simplest-known methods for each step: Boyd, Diaconis and Xiao [BDX04] provided a practical and simple subgradient method for computing $\lambda_{2}^{*}$ that can quite easily be implemented with GPU-acceleration in a modern library. Unfortunately, their method as-stated is a primal method only, and does not obtain a dual solution, so a slightly modified form would be necessary. The simplest dimension-reduction technique is that of Kwok, Lau, and Tung [KLT22], stated in Theorem 1.6. Unfortunately, the best-known bounds for $\alpha(G)$ (particularly for forbidden-minor graphs) are not currently as algorithmically simple or practical, nor is their application in Theorem 5.2 and Theorem 5.3. Finally, the simplest cut (rounding) algorithm is also likely that of Kwok, Lau, and Tung [KLT22, Section 3.2.3], although its constants could certainly use some improvement.
@@ Appendix A Graph Class Definitions
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus $g-1$ by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 6 for examples of these surfaces, and see Figure 7 for an example of a graph drawn on a genus-333 surface.
Figure 6: Examples of (left to right) a genus-111 surface, a genus-222 surface, a genus-333 surface, and a genus 101010 surface.
Figure 7: An example of a graph (left) and a drawing of the graph on a genus-333 surface (right).
@@
Definition A.3.
For graphs $G$ and $\hat{G}$, $\hat{G}$ is said to be a region intersection graph over $G$ if there is a mapping $f:V(\hat{G})\to 2^{V(G)}$ so that $G[f(v)]$ is connected for all $v\in V(\hat{G})$, and $uv\in V(\hat{G})$ if and only if $f(u)\cap f(v)\neq\emptyset$. We denote the set of region intersection graphs over $G$ as $\text{rig}(G)$. When $G$ is a planar graph, $\hat{G}\in\text{rig}(G)$ is said to be a string graph.
Figure 8: An example of strings in the plane (left), and their string graph (right).
A string graph is normally defined in a slightly different way, as the “geometric intersection graph” (a graph with vertices defined by geometric objects and edges defined by intersections) of “strings” (curves in the plane). See Figure 8 for an example of this construction. This is equivalent to a region intersection graph over a plane graph whose regions are specifically paths, which itself is equivalent to the above definition via some basic transformations.
The term “region intersection graph” in the literature usually refers to cases in which the “latent” graph $G$ has a forbidden $K_{h}$ minor, for some $h$. This generalizes the cases of bounded-genus and forbidden-minor graphs: If $G$ has no $K_{h}$ minor, and $\dot{G}$ is the barycentric subdivision of $G$ (which divides each edge with one additional vertex), then $G\in\text{rig}(\dot{G})$, and $\dot{G}$ also has no $K_{h}$ minor.
@@
A graph $G$ is said to be a $d$-dimensional $k$-ply neighbourhood system if it is the geometric intersection graph of a collection of $d$-dimensional balls $B_{1},\dots,B_{n}$ with ply at most $k$ for all points almost surely.
@@ Appendix B Vertex Expansion is NP-hard
We will show that edge expansion can be reduced to vertex expansion. Let $G=(V,E)$ be a graph with $n=|V|$ vertices, and let $\alpha>0$. These form the “input” to the edge expansion problem. We will show that there is a graph $G^{\prime}$ and a value $\beta$ so that $G$ has edge expansion at most $\alpha$ if and only if $G^{\prime}$ has vertex expansion at most $\beta$. At a high-level, $G^{\prime}$ will be formed by bisecting each edge of $G$, and then replacing each vertex that was not a bisection vertex with a clique of size $k$. For a sufficiently large value of $k$ (which is still polynomial in $n$), we will be able to show that an upper bound on the vertex expansion of $G^{\prime}$ can be “rounded” to an upper bound on the edge expansion of $G$. In particular, for any $\varepsilon>0$, we may assume $k$ to be a sufficiently large polynomial in $n$ so that for any $S\subset V$, $|S|k\leq|S|k+|E[S]|\leq|S|k+|E|\leq(|S|+\varepsilon)k$, We will place further requirements on $k$ (including the choice of $\varepsilon$) later.
We now give the explicit construction of $G^{\prime}$ and $\beta$, for a parameter $k$. where $E[S]$ denotes the set of edges in the induced subgraph of $G$ with the vertices $S$. Let $V_{1}$ be a set with $k$ copies of each $v\in V$ labelled $v^{1},\dots,v^{k}$, for a value $k$ to be chosen later. Let $V_{2}=E$. Let $E^{\prime}=\{ev^{i}:e\in E,v\in V\}$. Let $G^{\prime}$ be a graph with vertices $V^{\prime}=V_{1}\cup V_{2}$ and edges $E^{\prime}$. We choose $\beta:=\frac{\alpha}{k}$.
@@
We first claim that if there is a subset $S\subset V$, then there is a subset $S^{\prime}\subset V^{\prime}$ with $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$, where $E[S]$ denotes the edges in the induced subgraph $G[S]$. In particular, $S^{\prime}$ is exactly the subset of $V^{\prime}$ formed by $E[S]\subset V_{2}$ and $\{v^{i}:v\in S\}\subset V_{1}$.
@@
We next claim that if there is a subset $S^{\prime}\subset V^{\prime}$, then there is a second subset $S^{\prime\prime}\subset V^{\prime}$ with with $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$, and a subset $S\subset V$ so that $|\delta(S)|=|\partial(S^{\prime})|$ and $|S|k+|E[S]|=|S^{\prime\prime}|$. Specifically, we choose $S^{\prime\prime}$ to be the set that contains, for each $v\in V$, every element of $\{v^{1},\dots,v^{k}\}$ if and only if $S$ contains any element of this set, and furthermore $S^{\prime\prime}$ contains $e=uv\in V_{2}$ if and only if $S$ contains some $u^{i}$ and $v^{j}$. Note that this corresponds exactly to the sets that can be mapped back to sets $S\subset S$ with the desired correspondence via the inverse of the transformation in the previous paragraph.
@@
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{2}$ in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. Henceforth, we may assume that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For the vertices $e=uv\in V_{2}$ not in $S^{\prime}$ but in $S^{\prime\prime}$, such vertices must have all their neighbours included in $S^{\prime\prime}$. Hence, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed 111 to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 111 to the numerator and 0 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
@@
Will will now finally prove that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta=\frac{\alpha}{k}$, In particular, we claim that, for every $\alpha\geq 0$, there exists $S\subset V$ so that $\frac{|\delta(S)|}{|S|}\leq\alpha$ in $G$ if and only if there exists $S^{\prime}$ so that $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$ in $G^{\prime}$. Note that there are only a polynomial number of possible values of $\frac{|\delta(S)|}{|S|}$ (specifically, there are at most $|E|\cdot|V|$ possible values), so we may assume $\alpha$ is one such value, and that $k$ is chosen so that $\alpha\cdot(1+\varepsilon)$ is strictly less than any larger such value.
@@
In the backward direction, assume $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Then, use the prior construction to find $S$ so that $\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\alpha k$. Recall that $|S|k+|E[S]|\leq(|S|+\varepsilon)k$, so $\frac{|\delta(S)|}{(|S|+\varepsilon)k}\leq\alpha k$ and thus $\frac{|\delta(S)|}{|S|}\leq\alpha(1+\varepsilon)$. Hence, by the earlier choice of $k$, we also get $\frac{|\delta(S)|}{|S|}\leq\alpha$.
@@ Appendix C Deferred Proof of Subdivision Lemma
We now prove Lemma 3.14 by adapting an argument of Kelner [Kel04, Kel06b, Kel06a]. The proof will be similar to, but not quite the same as, that of Lemma 3.11. The primary differences will be that the paths corresponding to edges will be randomly sampled, and that the “uniformity” is only approximate. It seems quite plausible that a more general form of Lemma 3.11 could be extracted from the below proof, but the most straightforward proof method using these random paths results in an extra factor of $\Delta$. This factor is perfectly fine for proving Lemma 3.14, but poses issues for a generalized form of Lemma 3.11.
@@
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree >6>6> 6 are added from a hexagonal subdivision, nor does any existing vertex increase have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree >6>6> 6 remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.
Kelner [Kel04, Kel06a, Proof of Lemma 5.2] is able to show that there exists random variables $\pi_{V}:V\to V^{\prime}$ and $\pi_{E}:E\to\{\text{paths through $H$}\}$ with the following properties:
@@
    For each $uv\in E$, $\pi_{E}(uv)$ is dependent only on $\pi_{V}(u)$ and $\pi_{V}(u)$.
@@
$\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{uv\in E}\sum_{u^{\prime},v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right).$
Let $\rho:=\max_{uv\in E}|\pi_{E}^{*}(uv)|$ be the maximum length (in terms of edges) of a sampled path, so $\rho\in\operatorname{\mathcal{O}}_{\Delta}(2^{k})$. Choose $f_{G}(v):=f_{H}(\pi_{V}^{*}(v))-\frac{1}{n}\sum_{x\in V}f_{H}(\pi_{V}^{*}(x))$, and $y_{G}(v):=2\rho\frac{\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{x\in V}|f_{G}(x)|^{2}}$.
We start by showing that the constraints are satisfied. First, note that $\sum_{v\in V}f_{G}(v)=\sum_{v\in V}f_{H}(v)-\sum_{v\in V}f_{H}(v)=\overline{0}$, so the normalization constraint is satisfied. Next, consider some edge $uv\in E$. By Cauchy-Schwarz and the triangle inequality (similar to the argument in Lemma 3.11),
$y_{G}(u)+y_{G}(v)\geq\frac{|f_{G}(u)-f_{G}(v)|^{2}}{\sum_{x\in V}|f_{G}(x)|^{2}},$
so the remaining constraints are satisfied.

### v4

@@ 7 Conclusion and Future Work
In order to obtain guarantees on the performance of reweighted spectral partitioning, we have provided bounds on $\gamma^{(n)}(G)$ for a number of graph classes. However, very few of them are tight, even up to constant factors. We pose the following open questions:
  • For a graph $G$ of $n$ vertices with genus $g$, is it true that $\gamma^{(n)}(G)\lesssim\frac{g}{n}$? A positive answer would also imply a conjecture of Spielman and Teng [ST96, ST07, Conjecture 1].
  • For a graph $G$ of $n$ vertices with no $K_{h}$-minor, is it true that $\gamma^{(n)}(G)\lesssim\frac{h^{2}}{n}$? If true, this would be an even stronger result than the first question.
  • For a graph $G$ of $n$ vertices with no $K_{h}$-minor, can it be shown that $\gamma^{(1)}(G)\lesssim\gamma^{(n)}(G)\log h$? If true, this would strengthen Theorem 1.4. Such a result would likely involve a slightly different construction than padded partitions.
We also ask if Algorithm 2 could be replaced by an algorithm that does not directly depend on the partition oracle, but still obtains the same guarantees. By analyzing the argument of Biswal, Lee, and Rao [BLR10, Theorem 4.4], it can be shown that it would suffice to obtain a constant-factor pseudo-approximation algorithm for the obnoxious $k$-median problem (specifically with $k=\Omega(n)$), which asks for a set $S$ of $k$ points in an approximate metric space $(X,d)$ (satisfying the triangle inequality up to a constant factor) maximizing $\sum_{x\in X}\min_{s\in S}d(x,s)$. The intuition for this problem is that proximity to some types of facilities may be undesirable (e.g., garbage dumps). Here, “pseudo-approximation” means that it would suffice to obtain a set of $k^{\prime}=\Omega(n)<k$ points that produces a constant-factor approximation for the best solution with $k$ points. This problem is well-studied in practice [LC21, CMD18, Gok20, LG18, BLMN07, SCC25], but has seen only limited study in theory [KD22].
@@ 8 Acknowledgements
Thank you also to Lap Chi Lau for his excellent course in 2022, in which the some of the preliminary results of this work began as a course project.
In parallel to this work, and also as a part of the same course, Kam Chuen Tung independently derived essentially the same result as Proposition 5.5, and similar bounds for $\gamma^{(1)}$ based on $L^{2}$-extremal spread for bounded-genus graphs and $K_{h}$-minor-free graphs. These results have appeared in his recent PhD thesis [Tun25, Chapter 7], which also includes some interesting generalizations.
After the initial preprint of this work, Lap Chi Lau notified the author that he was also independently aware of the bound on $\gamma^{(1)}$ for planar graphs stated in Proposition 5.5 prior to either work.
@@ Appendix A Graph Class Definitions
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus g1g-1 by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 11 for examples of these surfaces, and see Figure 12 for an example of a graph drawn on a genus-$3$ surface.
Figure 11: Examples of (left to right) a genus-$1$ surface, a genus-$2$ surface, a genus-$3$ surface, and a genus $10$ surface.
Figure 12: An example of a graph (left) and a drawing of the graph on a genus-$3$ surface (right).
@@
A graph $G$ is said to be the intersection graph of $d$-dimensional balls with ply $k$ (sometimes called a $d$-dimensional $k$-ply neighbourhood system) if there is a collection of $d$-dimensional balls $B_{1},\dots,B_{n}$ with ply at most $k$ for all points almost surely, so that each ball corresponds to a vertex, and each edge corresponds to a pair of vertices whose balls intersect.
@@ Appendix B Vertex Expansion is NP-hard
We will show that edge expansion can be reduced to vertex expansion. Let $G=(V,E)$ be a graph with $n=|V|$ vertices, and let $\alpha>0$. These form the input to the edge expansion problem, and hence to our reduction. We will construct a graph $G^{\prime}$ and a value $\beta$ so that $G$ has edge expansion at most $\alpha$ if and only if $G^{\prime}$ has vertex expansion at most $\beta$. At a high-level, $G^{\prime}$ will be formed by bisecting each edge of $G$, and then replacing each vertex that was not a bisection vertex with a clique of size $k$. For a sufficiently large value of $k$ (which is still polynomial in $n$), we will be able to show that an upper bound on the vertex expansion of $G^{\prime}$ can be “rounded” to an upper bound on the edge expansion of $G$. In particular, for any $\varepsilon>0$, we may assume $k$ to be a sufficiently large polynomial in $n$ so that for any $S\subset V$, $|S|k\leq|S|k+|E[S]|\leq|S|k+|E|\leq(|S|+\varepsilon)k$, We will place further requirements on our choice of $k$ (including the choice of $\varepsilon$) later.
We now give the explicit construction of $G^{\prime}$ and $\beta$, for a parameter $k$. Let $E[S]$ denote the set of edges in the induced subgraph of $G$ with the vertices $S$. Let $V_{1}$ be a set with $k$ copies of each $v\in V$ labelled $v^{1},\dots,v^{k}$. Let $V_{2}=E$. Let $E_{1}=\{v^{i}v^{j}:v\in V,1\leq i,j\leq k\}$ and $E_{2}=\{ev^{i}:e\in E,v\in V,1\leq i\leq k\}$. Let $G^{\prime}$ be a graph with vertices $V^{\prime}=V_{1}\cup V_{2}$ and edges $E^{\prime}:=E_{1}\cup E_{2}$. We choose $\beta:=\frac{\alpha}{k}$.
We will henceforth use the notation $\partial A:=\{N_{G^{\prime}}(A)\cap V^{\prime}\setminus A\}$ for a set of vertices $A\subset V^{\prime}$, and the notation $\delta B:=E\cap(B\times(V\setminus B))$ for a set of edges $B\subset E$.
@@
We first claim that, if there is a subset $S\subset V$, then there is a subset $S^{\prime}\subset V^{\prime}$ so that $|\partial(S^{\prime})|=|\delta(S)|$ and $|S^{\prime}|=|S|k+|E[S]|$, where $E[S]$ denotes the edges in the induced subgraph $G[S]$. In particular, if we let $S^{\prime}$ be the subset of $V^{\prime}$ formed by the union of $E[S]\subset V_{2}$ and $\{v^{i}:v\in S,1\leq i\leq k\}\subset V_{1}$, then we actually have the stronger condition that $\partial(S^{\prime})=\delta(S)$.
@@
We next claim that for any subset $S^{\prime}\subset V^{\prime}$, there are subsets $S^{\prime\prime}\subset V^{\prime}$ and $S\subset V$ so that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$, $|\delta(S)|=|\partial(S^{\prime})|$ and $|S|k+|E[S]|=|S^{\prime\prime}|$. Specifically, we choose $S^{\prime\prime}$ to be the set that contains, for each $v\in V$, every element of $\{v^{1},\dots,v^{k}\}$ if and only if $S$ contains any element of this set, and furthermore $S^{\prime\prime}$ contains $e=uv\in V_{2}$ if and only if $S$ contains some $u^{i}$ and $v^{j}$. Note that this choice of $S^{\prime\prime}$ corresponds to a set that can be mapped back to sets $S\subset V$ with the desired correspondence via the inverse of the transformation in the previous paragraph.
@@
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{1}$ that are in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least $1$ to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and $1$ to the denominator. Henceforth, we may assume without loss of generality that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For each vertex $e=uv\in V_{2}=E$ in $S^{\prime\prime}$ but not in $S^{\prime}$, its neighbours $N_{G^{\prime}}(e)$ are a subset of $S^{\prime\prime}$. Hence, it contributed at least $1$ to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and $1$ to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed $1$ to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute $1$ to the numerator and 0 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
@@
We will now finally prove that $\phi(G)\leq\alpha$ if and only if $\psi(G^{\prime})\leq\beta=\frac{\alpha}{k}$. In particular, we claim that, for every $\alpha\geq 0$, there exists $S\subset V$ so that $\frac{|\delta(S)|}{|S|}\leq\alpha$ in $G$ if and only if there exists $S^{\prime}$ so that $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$ in $G^{\prime}$. Note that there are only a polynomial number of possible values of $\frac{|\delta(S)|}{|S|}$ (specifically, there are at most $|E|\cdot|V|$ possible values), so we may assume $\alpha$ is one such value, and that $k$ is chosen so that $\alpha\cdot(1+\varepsilon)$ is strictly less than any larger such value.
@@
In the backward direction, assume $\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Then, use the prior construction to find $S$ so that $\frac{|\delta(S)|}{|S|k+|E[S]|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}\leq\frac{\alpha}{k}$. Recall that $|S|k+|E[S]|\leq(|S|+\varepsilon)k$, so $\frac{|\delta(S)|}{(|S|+\varepsilon)k}\leq\frac{\alpha}{k}$ and thus $\frac{|\delta(S)|}{|S|}\leq\alpha(1+\varepsilon)$. Hence, by the earlier choice of $k$, we also get $\frac{|\delta(S)|}{|S|}\leq\alpha$.
@@ Appendix C Deferred Proof of Subdivision Lemma
We now prove Lemma 5.17 by adapting an argument of Kelner [Kel04, Kel06b, Kel06a].
Lemma C.1 (Restatement of Lemma 5.17).
The proof will be similar to, but not quite the same as, that of Lemma 5.14. The primary differences will be that the paths corresponding to edges will be randomly sampled, and that the “uniformity” is only approximate. It seems quite plausible that a more general form of Lemma 5.14 could be extracted from the below proof, but the most straightforward proof method using these random paths results in an extra factor of $\Delta$. This factor is perfectly fine for proving Lemma 5.17, but poses issues for a generalized form of Lemma 5.14, so we have kept this proof and construction separate.
@@
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree >6>6 are added from a hexagonal subdivision, nor does any existing vertex have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree >6>6 remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.
Kelner [Kel04, Kel06a, Proof of Lemma 5.2] is able to show that there exist random variables $\pi_{V}:V\to V^{\prime}$ and $\pi_{E}:E\to\{\text{paths through $H$}\}$ with the following properties:
@@
    For each $uv\in E$, $\pi_{E}(uv)$ is dependent only on $\pi_{V}(u)$ and $\pi_{V}(v)$.
@@
$\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{uv\in E}\sum_{u^{\prime},v^{\prime}\in p(uv)}y_{H}(u^{\prime})+y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right).$
Let $\rho:=\max_{uv\in E}|\pi_{E}^{*}(uv)|$ be the maximum length (in terms of edges) of a sampled path, so $\rho\in\operatorname{\mathcal{O}}_{\Delta}(2^{k})$. Choose $f_{G}(v):=f_{H}(\pi_{V}^{*}(v))-\frac{1}{n}\sum_{x\in V}f_{H}(\pi_{V}^{*}(x))$, and $y_{G}(v):=2\rho\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(x)$.
We start by showing that the constraints are satisfied. By definition, $\sum_{v\in V}f_{G}(v)=\overline{0}$. Consider some edge $uv\in E$, and the corresponding path $\pi_{E}^{*}(uv)$ through $H$. By Cauchy-Schwarz and the triangle inequality (similar to the argument in Lemma 4.5),
$y_{G}(u)+y_{G}(v)\geq|f_{G}(u)-f_{G}(v)|^{2},$
so the constraints are satisfied.